# A new perspective on modulus stabilization

### v2

@@ VI Criteria for bulk potential to ensure modulus stabilization
We did a quick check for the case with no bulk interaction i.e. $V(\Phi)=\frac{1}{2}m^{2}\Phi^{2}$ with a set of parameters for which modulus stabilization is possible (see Eq.(2.22), [5]): $\alpha=0.2,k=1,v=0.05,\epsilon=\frac{m^{2}}{4k^{2}}=0.5$. This satisfies Eqs.(70) and (73) with $r_{c}^{*}=1.156$. Similarly, one can do the checking for other potentials but these relations are guaranteed to hold as they are derived generally.
VII Conclusion
Our analysis brings out a deep correlation between the stability of the scalar field (which has been employed for modulus stabilization) and the corresponding existence of a stable minimum for the modulus. The outcomes can be summarized as follows: (1) Existence of a global minimum for the scalar sector immediately ensures a stable minimum for the modulus. (2) On the contrary, the scalar sector which does not have a global minimum may or may not lead to modulus stabilization. It is to be noted that the results are subject to the two types of boundary conditions that we assumed. (3) Our calculations addressed various scalar field actions to explore the resulting stabilization of the moduli sector. This included a canonical kinetic term as well as phantom like term for the scalar sector. Moreover, choosing different forms of the scalar field potentials, we carefully examined the possibility of a corresponding modulus stabilization. (4) It may further be observed that such different stabilizing scalar sectors in the scalar-tensor models have direct correspondence with higher curvature gravity models and thus, our calculations indirectly include various viable f(R) models which may lead to modulus stabilization. The results therefore provide us with a powerful tool to test modulus stabilization for different bulk stabilizing fields.
Acknowledgements.
SB is supported through INSPIRE-SHE Scholarship by the Department of Science and Technology (DST), Government of India.
@@ Appendix A Including backreaction - The Superpotential Method
In GW scenarios, we neglected the back-reaction of the metric to the presence of the scalar field in the bulk but this is important indeed and therefore, it would be very nice to simultaneously solve the Einstein and the bulk scalar equations, to have the back-reaction exactly under control. DeWolfe et al. [4] provides us with such a formalism. Denote the scalar field in the bulk by $\Phi$, and consider the action
$\begin{split}\int d^{5}x\sqrt{g}\left[-M^{3}R++\frac{1}{2}(\nabla\Phi)^{2}-V(\Phi)\right]\\ -\int d^{4}x\sqrt{g_{4}}\lambda_{hid}(\Phi)-\int d^{4}x\sqrt{g_{4}}\lambda_{vis}(\Phi)\end{split}$ (74)
@@
The bulk equations Eqs. (76-78) along with the boundary conditions Eq. (A) form the coupled gravity-scalar system. We can now define the function $W(\Phi)$ via the equations
@@ Appendix B A relation between $r_{c}$ and p for the BFG case
Given that we take our conjecture holds true that there is indeed modulus stabilization for all BFG type potentials, a natural question to ask is whether we can establish some relation between $r_{c}$ and the only parameter in the problem p. We listed down all known $r_{c}$ values till p = 19, plotted $r_{c}$ vs p and found the empirical expression which best-fits the curve as

### v4

@@ 6 General mathematical conditions for the bulk potential to ensure modulus stabilization
We did a quick check for the case with no bulk interaction i.e. $V(\Phi)=\frac{1}{2}m^{2}\Phi^{2}$ with a set of parameters for which modulus stabilization is possible (see Eq.(2.22), Chacko1): $\alpha=0.2,k=1,v=0.05,\epsilon=\frac{m^{2}}{4k^{2}}=0.5$. This satisfies Eqs.(66) and (69) with $r_{c}^{*}=1.156$. Similarly, one can do the checking for other potentials but these relations are guaranteed to hold as they are derived generally.
The radion mass squared is given by Eq. (67) times $\frac{e^{2k\pi r_{c}^{*}}}{(Ak\pi)^{2}}$, since the canonical radion field is defined as $\varphi\equiv Ae^{-k\pi r(x)}$, where $A=\sqrt{24M^{3}/k}$. Thus, an overall warping of $e^{-k\pi r_{c}^{*}}$ emerges for the radion mass and it is expected to be $\mathcal{O}(TeV)$ in agreement with the existing literature GW2; csaki2000cosmology; csaki2001. If one focuses on the positive terms in Eq. (6.4), then we can say that if a bulk potential is steep in the sense that $V^{\prime\prime}(\Phi_{OR})$ is large, and admits a rapidly varying analytic solution $\Phi_{OR}(\pi r_{c})$ with $r_{c}$, atleast at $r_{c}^{*}$, then it ensures a heavy radion. The fields living on the visible brane couple to the radion through the induced metric kofman2004exact; csaki2001, with an interaction $\frac{\varphi}{\sqrt{6}\Lambda_{W}}T^{\mu}_{\mu}$ where $\Lambda_{W}=M_{Pl}e^{-kr_{c}^{*}}\sim O({\rm TeV})$, $M^{2}_{Pl}=(1-e^{-2kr_{c}^{*}})/(k\kappa^{2})\sim 1/(k\kappa^{2})$, $T^{\mu}_{\mu}$ is the trace of the physical energy-momentum tensor of the TeV brane fields and $\kappa^{2}$ is the 5D Newton’s constant. It is then clear that the radion couples as $\sim 1/$TeV to the Standard Model fields. An acceptable phenomenology requires $\mathcal{O}(TeV)$ radion mass which, as we saw, is likely in this context. As the dependence of the coupling on bulk potential is through the value of the stabilized modulus, it can be understood by analyzing Eq. (66) for a given bulk potential, but a general feature is difficult to infer. Bounds on the radion mass and its coupling to visible sector fields can be found in Ref. bae2001bounds.
7 Stabilized modulus from brane world sum rules?
An essential feature one may have noticed while reading the paper till this point is that the analytic (or even numeric) calculation of the stabilized radion (if it gets stabilized) in each of the cases has been quite laborious. This in fact can be seen in several papers dealing with radion physics. So, is there an alternative way by which we could arrive at the stabilized modulus value - even in some cases? This is the question we want to address in this section.
Gibbons, Kallosh and Linde Gibbons put down a set of consistency conditions for several braneworld scenarios, derived purely from the Einstein equations. These are widely known as the brane world sum rules or the GKL sum rules (after the authors). An extension of these rules in arbitrary dimensions was later provided by Leblond et al. Leblond. [^8] We refer the reader to Ref. Leblond for a quick and self-contained review of the brane world sum rules - we do not go over the derivation in this paper. In this section, we show for the first time, that in certain Goldberger-Wise scenarios where stabilization occurs, the GKL rules can potentially help pin down the value of the stabilized modulus. This holds true even when gravitational contribution to the radion potential is included.
One of the GKL conditions is the check that the summation of the flat brane tensions and the non-negative integral of the gradient energy of the bulk scalars must vanish. The original GW scheme did not respect this criterion and hence was the need to involve backreaction and solve for the combined gravity-scalar system. However, in generalized GW scenarios (see Sec.3), the choice of a linear visible brane potential and the fine-tuning of the 4D cosmological constant to be negligible lets us use the brane sum rules to get to a non-trivial inequality expression involving the stabilized modulus ($r_{c}$) and the parameters in our setup. [^9] In both the cases discussed below, the parameters $\alpha$ and $v$ are taken to be small and almost of the same order.
The brane world sum rule of interest to us is thus given by (see Section 2, Gibbons)
$\sum_{i=\{vis,hid\}}\lambda_{i}(\Phi_{i})+\oint\Phi^{\prime}\cdot\Phi^{\prime}=0\ .$ (70)
where $\lambda_{i}(\Phi_{i})$ is the flat brane tension (including the brane potential for the scalar field $\Phi$) on the corresponding (visible or hidden) brane, and the second term corresponds to the non-negative gradient energy of the scalar. As a result, the brane tensions must obey
$\sum_{i=\{vis,hid\}}\lambda_{i}(\Phi_{i})\leq 0$ (71)
As per our setting for the generalized GW scenarios, the $\lambda_{i}$s take the form
$\lambda_{hid}=\gamma(\Phi(0)^{2}-k^{3}v^{2})^{2}+T_{hid}$ (72)
$\lambda_{vis}=2k^{5/2}\alpha\>\Phi(\pi)+T_{vis}$ (73)
where $T_{hid}$ and $T_{vis}$ are the brane tensions in absence of the scalar field. Recalling $\Lambda_{4D}=(T_{hid}+\frac{\Lambda_{b}}{k})/k^{4}$ to be the 4D cosmological constant which we tuned to be small as in Ref. Chacko1, $T_{hid}\approx-\Lambda_{b}/k$. Therefore, $\tau=(T_{vis}+T_{hid})/k^{4}$. A thing to notice is that in the original GW scheme, $\Phi(0)$ was not exactly equal to $k^{3/2}v$. Rather, they assumed a large coupling constant ($\gamma$) using which one could show that $\gamma(\Phi(0)^{2}-k^{3}v^{2})^{2}=\frac{m^{4}}{64k^{2}\gamma}$. This vanished in the limit $\gamma\rightarrow\infty$. However, in our present setup, we assume the boundary condition $\Phi(0)=k^{3/2}v$ which straightaway sets the first term in Eq.(72) to zero.
It might be confusing to see that $(T_{hid}+T_{vis})$ does not vanish. This comes because we promote the flat 4D background to a general metric. As a general metric no longer requires the fine-tuning in Eq. (6), the $(T_{hid}+T_{vis})$ sum is non-zero and thereby ensures a non-zero gravitational contribution to the radion potential (see Sec.2.2). When we do not replace $\eta_{\mu\nu}$ with $g_{\mu\nu}$, the fine-tuning remains intact and the gravitational contribution vanishes as has mostly been the case in this paper till now.
Adding Eqs.(72) and (73), we get
$\lambda_{hid}+\lambda_{vis}=2k^{5/2}\alpha\>\Phi(\pi)+k^{4}\tau$ (74)
Finally, plugging this result into Eq.(71) gives us
$2\alpha\>\Phi(\pi)+k^{3/2}\tau\leq 0$ (75)
In what follows, we will work with the above criterion which serves as a tailored brane sum rule for such generalized GW scenarios. The only unknown in Eq.(75) which is determined by our choice of the bulk scalar potential is the value of the scalar field at the visible brane i.e. $\Phi(\pi)$. Hence, the only step needed to get the inequality expression for the stabilized modulus is to solve for the scalar field profile.
As discussed previously, a class of bulk scalar potentials where radion stabilization is guaranteed is where they possess a global minimum. Thus, we consider a scalar field where the bulk potential is dominated by the mass term,
$V_{b}(\Phi)=\frac{1}{2}m^{2}\Phi^{2}\;$ (76)
Following the steps as in Sec.3, one then obtains the following solution for the scalar field profile
$\Phi(\phi)=-\frac{k^{3/2}\alpha}{4}e^{4kr_{c}(\phi-\pi)}+k^{3/2}ve^{-\epsilon kr_{c}\phi}\;$ (77)
where $\epsilon=m^{2}/4k^{2}$ with $|\epsilon|\ll 1$ (as $\alpha v>0$, $m^{2}$ has to be positive to be consistent with the expression for mass of the radion Chacko1). Now, following the steps as discussed in Sec.3.2 allows one to get to the expression for the radion potential up to the leading order in $v$ and $\epsilon$, including the gravitational contributions Chacko1.
$V(\varphi)=2k^{4}\alpha v\left(\frac{\varphi}{A}\right)^{4+\epsilon}+k^{4}\tau\left(\frac{\varphi}{A}\right)^{4}\;$ (78)
where $\tau$ is redefined as $\tau\coloneq(\tau-\frac{\alpha^{2}}{4})$. The minimum of this potential occurs for the condition:
$e^{-\epsilon kr_{c}\pi}=-\frac{\tau}{2\alpha v}+\mathcal{O}(\epsilon)$ (79)
Thus, we see that Eq.(79) gives us an expression for the stabilized modulus in this setup. It should be noted that $\alpha v>0$ enforces $\tau$ to be negative, the left side of the equation being an exponential. Now, lets turn our attention to the brane world sum rules and see if we could get somewhere close to this result. Plugging in the solution $\Phi(\pi)$ from Eq.(77) into the tailored brane sum rule for such cases (Eq.75) gives us
$\tau-\frac{\alpha^{2}}{4}+2\alpha ve^{-\epsilon kr_{c}\pi}\leq 0$ (80)
This leads us to the following inequality expression for the stabilized modulus ($r_{c}$),
$e^{-\epsilon kr_{c}\pi}\leq-\frac{\tau}{2\alpha v}+\frac{\alpha}{8v}$ (81)
As both $\alpha$ and $v$ are small parameters and almost of the same order, we can drop the second term in Eq.(81) in favour of the first term which is indeed dominating. This allows us to get to the result
$e^{-\epsilon kr_{c}\pi}\leq-\frac{\tau}{2\alpha v}$ (82)
Thus, we get an effective lower bound on $r_{c}$ and we notice that this very well coincides with the true value of the stabilized modulus. If the gravitational part were zero, then the right side of Eq. (81) becomes $\frac{\alpha}{4v}$. As the stabilized radion value from Eq. (79) would then be $\frac{\alpha}{8v}$, we see that the non-trivial bound gives the correct order of magnitude.
Can the brane-sum bounds be helpful in other scenarios? - say for the bulk cubic interaction. This calculation was already performed in Sec.4.2 (including gravitational contributions, the $\frac{\alpha^{2}}{4}$ in Eq.(30) and further equations should be replaced by $\tau$ - the approximate stabilized modulus value will not change). Thus, plugging in the solution of the scalar field from Eq.(28) into Eq.(75) gives us
$\frac{4\alpha v}{1+\xi kr_{c}\pi}+2\tau-\alpha^{2}\leq 0$ (83)
Assuming $\frac{4\alpha v}{1+\xi kr_{c}\pi}$ and $\alpha^{2}-2\tau$ have the same sign  [^10] If they were taken to be of opposite signs, the nature of bound in Eq.(84) and consequently Eq.(85) would be flipped i.e. a lower bound on $r_{c}$., we get the following bounding criterion for the stabilized modulus,
$k\pi r_{c}\leq-\frac{1}{\xi}+\frac{4\alpha v}{\xi(\alpha^{2}-2\tau)}$ (84)
As $\xi\small\propto v$, the second term is independent of $v$ and hence, is proportional to $\frac{1}{\eta(\alpha-2\tau/\alpha)}$. Now, $\alpha$ being a small parameter and $\eta$ being not so small (as the cubic term dominates the bulk potential) make the denominator very large, thus making the second term contribute negligibly compared to the first one. Therefore,
$k\pi r_{c}\leq-\frac{1}{\xi}$ (85)
Hence, we get an upper bound on $r_{c}$ which due to its proximity to the true value (Eq.32) is yet again, a great estimate for the stabilized modulus.
8 Discussion
The study brings out a deep correlation between the form of the bulk scalar potential - more generally, the scalar field profile (which has been employed for modulus stabilization) and the corresponding existence of a stable minimum for the radion potential. Our calculations addressed various scalar field actions to explore the resulting stabilization of the moduli sector. This included a canonical kinetic term as well as phantom like term for the scalar sector. Moreover, choosing different forms of the scalar field potentials, we carefully examined the possibility of a corresponding modulus stabilization. A summary of these cases can be found in Sec.4.7. We also demonstrated that the qualitative nature of bulk potentials can hint at the fate of modulus stabilization. In particular, the existence of a global minimum for the scalar sector was shown to ensure a stable minimum for the modulus. More precisely, the existence of a global minimum in the bulk potential is sufficient but not a necessary condition for modulus stabilization. On the contrary, the scalar sector which does not have a global minimum may or may not lead to radion stabilization. Taking into account the stability of the bulk scalar, this situation gets further classified: a stable bulk scalar always stabilizes the modulus, and failing to stabilize the radion can only result from an unstable bulk scalar. We should remember that all these conclusions hold provided an appropriate extremum for the radion potential exists in the first place. We then deduced the necessary criteria that any bulk scalar potential must satisfy to enable radion stabilization. It must be noted that our analysis consistently employed the CMS boundary conditions. The physical motivation behind choosing these conditions was explained in Sec.3.1. In principle, one is free to explore innumerable boundary conditions - each of these may have its own physical justification and they all serve as potential future works. The only instance where we deviated from CMS boundary conditions was when we discussed a non-trivial backreacted geometry with the BFG potentials. In this regard, we should mention that even under the original Goldberger-Wise boundary conditions, a global minimum in bulk potential stabilized the radion, whereas potentials with a global maximum and no global minimum (both the stable and unstable bulk scalars) failed to stabilize the radion. It may further be observed that such different stabilizing scalar sectors in the scalar-tensor models have direct correspondence with higher curvature gravity models and thus, our calculations indirectly include various viable $f(R)$ models of gravity which may lead to modulus stabilization banerjeepaul; shafaq1.
Regarding the CMS boundary conditions, we already highlighted their physical motivation, but one must also remember that they are most naturally motivated by our singular perturbation theory-based solution technique. If one sees our outer region (OR) and boundary region (BR) equations (see Sec.3.1), we certainly need a Dirichlet boundary condition at the hidden brane since there is only a first order derivative of $\Phi$ in the OR equation, whereas the BR equation demands a Neumann boundary condition at the visible brane (since there is a second order derivative of $\Phi$ in the BR equation). This is ensured by a linear visible brane potential. Hence, we are naturally led to the CMS boundary conditions to make the solution technique work. If one wishes to work with the original GW Dirichlet boundary conditions on both branes, then the present analytic solution technique and they have to resort to numerics. Even then, they have to check large parameter spaces before reaching any general result.
While the five-dimensional Randall–Sundrum (RS1) model provides an elegant resolution to the gauge hierarchy problem, it faces increasing tension with present collider constraints. In particular, the non-observation of an $\mathcal{O}(\text{TeV})$ Kaluza–Klein graviton mode at the ATLAS and CMS experiments has progressively restricted the viable parameter space of the model CMS:2024nht. To remain consistent with these bounds, one must now invoke a modest hierarchy of order $\lesssim 10^{-2}$ between the five-dimensional Higgs vacuum expectation value and the five-dimensional Planck scale—an assumption that introduces its own theoretical challenges. In particular, the fine-tuning problem appears in a new form and spoils the elegance of the original RS1 model, whose success was to resolve the hierarchy problem of the Higgs mass without fine-tuning of the fundamental parameters (see, e.g., the introductory discussions of arun2015graviton; bhaumik2024moduli). A natural way to circumvent these issues is to consider higher-dimensional generalizations of the RS framework with multiple warping, which have been extensively studied in Refs. livingedge; mukhopadhyaya2012matter; chakraborty2014bulk; arun2015graviton; bhaumik2022moduli; bhaumik2023nested; bhaumik2024moduli. These multiply warped models can remain compatible with current collider bounds while continuing to address the hierarchy problem effectively. It would therefore be worthwhile to extend and apply the present methodology to such higher-dimensional setups in future works.
This paper also discusses a potential connection between geometric consistency conditions and modulus stabilization. We showed that in cases where the bulk scalar potential possesses a global minimum and thereby stabilizes the radion, the GKL sum rules can help us pin down the value of the stabilized radion. This may seem unexpected at first since getting to the stabilized modulus is different physics altogether. However, we could argue with a possible explanation - since there are an infinite number of GKL rules (they form a one-parameter family of conditions), these constraints essentially reduce the solution space of the modulus to a point that the only self-consistent solution is the stabilized one (subject to future exploration) - but still, why so? - we do not have an answer yet. Also, in the specific cases that we discussed, the fact that the visible brane potential did not vanish when evaluated at $\phi=\pi$ did the main trick as it gave away the radion contribution when plugged into the sum rule. Exploring other boundary conditions and brane potentials with such properties is thus quite lucrative. We also demonstrated that upon bulk cubic interaction, the sum rules offer the same insight. Even if one is hesitant to view this as an exact determination, it still serves as a great order estimation protocol. Thus, though this paper provides the first steps, whether the connection between brane sum rules and modulus stabilization goes far beyond the cases that we discussed, and whether we can understand any possible underlying physics are subject to further research. Some of these will be addressed in our future works.
Acknowledgements
SB acknowledges helpful discussions with Prof. Sourov Roy and Prof. Dilip Kumar Ghosh. We thank both the anonymous referees for their insightful comments which have led to the enhancement of this manuscript. SB is supported through INSPIRE-SHE Scholarship by the Department of Science and Technology (DST), Government of India.
Data Availability Statement
Being a theoretical study, no experimental data is associated with this work.
Code Availability Statement
Apart from numerical integration, no major code is associated with this work. This shall be made available on reasonable request.
@@ Appendix A Including backreaction - the superpotential method
In GW scenarios, we neglected the back-reaction of the metric to the presence of the scalar field in the bulk but this is important indeed and therefore, it would be very nice to simultaneously solve the Einstein and the bulk scalar equations, to have the back-reaction exactly under control. DeWolfe et al. Freedman provides us with such a formalism. Denote the scalar field in the bulk by $\Phi$, and consider the action
$\begin{split}\int d^{5}x\sqrt{g}\left[-M^{3}R++\frac{1}{2}(\nabla\Phi)^{2}-V(\Phi)\right]-\int d^{4}x\sqrt{g_{4}}\lambda_{h}(\Phi)-\int d^{4}x\sqrt{g_{4}}\lambda_{v}(\Phi)\end{split}$ (86)
@@
The bulk equations Eqs. (88-90) along with the boundary conditions Eq. (A) form the coupled gravity-scalar system. We can now define the function $W(\Phi)$ via the equations
@@
These jump conditions are satisfied only when the brane potentials are chosen of the form
$\lambda_{h}=W(\Phi_{h})+W^{\prime}(\Phi_{h})(\Phi-\Phi_{h})+\gamma_{h}(\Phi-\Phi_{h})^{2}$ (95)
and
$\lambda_{v}=-W(\Phi_{v})-W^{\prime}(\Phi_{v})(\Phi-\Phi_{v})+\gamma_{v}(\Phi-\Phi_{v})^{2}$ (96)
@@ Appendix B A relation between $r_{c}$ and p for the BFG case
Given that we take our conjecture holds true that there is indeed modulus stabilization for all BFG type potentials, a natural question to ask is whether we can establish some relation between $r_{c}$ and the only parameter in the problem p. We listed down all known $r_{c}$ values till p = 19, and found the empirical expression which best-fits the $r_{c}$ vs p curve as